The cost of the gold bracelet was $20.00 more than three times the silver one. Together they cost $320.00. How much did each cost?
Answer by checkley77(12569)

X+3X+20=320
4X+20=320
4X=320-20
4X=300
X=300/4
X=$75 IS THE COST OF THE SILVER BRACELET.
3*75+20=225+20=$245 FOR THE GOLD BRACELET.
PROOF:
75+245=320
320=320